It is desired to obtain the highest possible volume of 1.25 M urea, [Urea (aq)], from the following sources: 345 mL of 1.29 M urea solution, 485 mL of 0.653 M urea solution and 835 mL of 0.775 M urea solution. How can it be done? What is the maximum volume of solution that can be obtained?


1
Expert's answer
2021-08-28T06:26:34-0400

By calculating the moles of urea

Moles of 1st urea "=1.29×\frac{345}{1000}=0.45moles"


Moles of 2nd urea "=0.653×\frac{485}{1000}=0.31moles"


Moles of 3rd urea "=0.775×\frac{835}{1000}=0.65moles"